Two variables are defined, a regression equation is given, and one data point is given. Hgt = height in inches Age age in years of a child Hgt -24.2 +2.78(Age) The data point is a child 12 years old who is 60 inches tall. (a) Find the predicted value for the data point and compute the residual. Enter the exact answers. Predicted value = inches Residual inches
Answer #1

Predicted value = 24.2+2.78*12 = 57.56 inches

Residual = actual - predicted = 60 - 57.56 = 2.44 inches
Answer #2
Predicted Value- 57.28 Residual- 2.72